OpenStudy (anonymous):

15 is 375% of what number?

OpenStudy (amistre64):

part = % * total they give you part and %

OpenStudy (amistre64):

total = 15/3.75

OpenStudy (anonymous):

375 = 15/n n=15/375

OpenStudy (anonymous):

opps my bad, forgot 375 = 15/n * 100, it should be 15/3.75